# Large Scale Knowledge Washing

### v2

@@ Abstract
Large language models show impressive abilities in memorizing world knowledge, which leads to concerns regarding memorization of private information, toxic or sensitive knowledge, and copyrighted content. We introduce the problem of Large Scale Knowledge Washing, focusing on unlearning an extensive amount of factual knowledge. Previous unlearning methods usually define the reverse loss and update the model via backpropagation, which may affect the model’s fluency and reasoning ability or even destroy the model due to extensive training with the reverse loss. Existing works introduce additional data from downstream tasks to prevent the model from losing capabilities, which requires downstream task awareness. Controlling the tradeoff of unlearning and maintaining existing capabilities is also challenging. To this end, we propose LaW (Large Scale Washing) to update the MLP layers in decoder-only large language models to perform knowledge washing, as inspired by model editing methods and based on the hypothesis that knowledge and reasoning are disentanglable. We derive a new objective with the knowledge to be unlearned to update the weights of certain MLP layers. Experimental results demonstrate the effectiveness of LaW in forgetting target knowledge while maintaining reasoning ability. The code will be open-sourced at https://github.com/wangyu-ustc/LargeScaleWashing.
@@ 1 Introduction
Large Language Models (LLMs) are shown to memorize extensive knowledge or factual relations (Chen et al., 2022; Alivanistos et al., 2022; Youssef et al., 2023). However, the memorization of knowledge in LLMS raises both moral and legal concerns. Factual knowledge may involve personal and sensitive information whose memorization can violate strict regulations (Legislature, 2018; Act, 1996; Parliament and of the European Union, 2016), and memorizing copyright content is also problematic – The New York Times [^1] https://nytco-assets.nytimes.com/2023/12/NYT_Complaint_Dec2023.pdf recently filed lawsuit against OpenAI to protect its copyright of articles.
@@
To prevent the undesired memorization of the above-mentioned knowledge, the simplest solution is perhaps to label data that has the potential to raise concerns in advance and exclude sensitive data from the pre-training stage. However, this solution needs exhaustive human effort and may not be feasible as the pretraining corpus for LLM is normally extremely large. This impracticality motivates the study of machine unlearning (Liu et al., 2024a; Yao et al., 2024; Si et al., 2023; Yao et al., 2023b; Zhang et al., 2023). When there are concerns about memorizing sensitive knowledge, these methods aim to update the LLM to forget that knowledge with a relatively small computational cost. Most of these methods are in the paradigm of defining an “unlearning" loss (essentially the reverse loss of Next-Word-Prediction on the unlearning dataset) and updating the full models by backpropagating from the loss. However, updating the model with backpropagation may hurt the model’s abilities in downstream tasks requiring reasoning. When the knowledge to be unlearned scales up, it may require extensive updates of the model parameters, which could even destroy the model (as shown in our experiments). Some efforts to overcome this limitation define a “utility" loss from specific downstream tasks and optimize both unlearning and utility losses (Liu et al., 2024a). However, the applications of these methods may be limited when we focus on the generalizability of LLMs where no downstream tasks are specified.
In this work, we focus on the Large Scale Knowledge Washing problem: how do we unlearn the knowledge at scale (termed knowledge washing) from large language models without affecting the model’s reasoning ability? (as shown in Figure 1) We hypothesize that the knowledge and reasoning abilities in LLMs are disentanglable, which gives rise to a feasible solution to the above problem. To do so, we design a novel method named LaW (Large Scale Washing), which is inspired by model editing methods (Meng et al., 2022, 2023), where MEMIT (Meng et al., 2023) can perform massive knowledge editing by identifying a subset of parameters in LLMs that account for the prediction of certain knowledge and then editing these parameters with a closed-form equation. In the context of knowledge washing, we follow this idea to first identify a subset of parameters, and then derive a new objective to update the parameters. In the knowledge-washing context, we can have a similar closed-form solution as in the model-editing scenario. However, some important variables in this solution are not available. This leads to our new objective, which requires optimization rather than having a closed-form solution. Thus we introduce some practical considerations to facilitate the optimization. The key difference between LaW and model-editing ideas lies in the fact that model-editing methods propose to add factual relations in the model weights while LaW aims to delete factual relations.
We evaluate our methods using two small-scale datasets and a newly created large-scale dataset derived from Wikipedia triplets, which includes 332,036 facts. Across different methods, we compare the accuracy and QA-F1 scores of the generated answers from prompts formed from the triplets, demonstrating that our approach achieves the most thorough knowledge washing. Importantly, LaW maintains the model’s reasoning abilities, as verified by its performance on various reasoning tasks.
@@ 2 Related Work
Recent research has increasingly focused on the concept of machine unlearning in the context of large language models (LLMs), highlighting both its challenges and necessities (Liu et al., 2024a; Yao et al., 2024; Si et al., 2023; Yao et al., 2023b; Zhang et al., 2023). Beyond addressing privacy concerns necessitating unlearning in LLMs, several studies have employed unlearning techniques to investigate the influence of specific subsets of training data on model performance (Isonuma and Titov, 2024; Zhao et al., 2024). To facilitate knowledge unlearning, various approaches have been proposed. One method involves retraining the LLM on the targeted dataset using a reverse loss function, coupled with training on an irrelevant dataset to preserve performance on unrelated tasks. This can be implemented through the addition of unlearning layers (Chen and Yang, 2023) or directly within the large language model itself (Eldan and Russinovich, 2023). Unlike these approaches, which apply to whole sequences in the unlearning subset, Wang et al. (2024) suggest focusing on specific spans within sequences to minimize disruption to unrelated tasks (Wang et al., 2024). Furthermore, an alternative strategy known as in-context unlearning utilizes few-shot prompts to induce forgetting of specific datasets directly within the context of use, presenting a different approach from traditional training-based methods (Pawelczyk et al., 2023). In a distinct line of research, other methods target the mitigation of harmful outputs by collecting problematic prompts and applying techniques such as instruction tuning (Liu et al., 2024b) or reinforced learning (Lu et al., 2022) to prevent toxic responses.
Model Editing of LLMs
@@
Model editing in large language models pertains to the modification of factual relations within the models to integrate new world knowledge (Yao et al., 2023a). Initial approaches to model editing focused on single-fact adjustments, requiring the model to update one factual relation at a time. Prominent methods in this domain include ROME (Meng et al., 2022), MEND (Mitchell et al., 2022a), T-Patcher (Huang et al., 2023), and IKE (Zheng et al., 2023). These techniques, however, often face stability issues after multiple edits, complicating the process of batch editing, where multiple new factual relations are introduced simultaneously. In response to these challenges, advanced methods like GRACE (Hartvigsen et al., 2022) and SERAC (Mitchell et al., 2022b) have been developed for effective batch editing. Further advancements have tested these methodologies on larger models, such as GPT2-XL and GPT-J-6B, with techniques like MEMIT (Meng et al., 2023) and Model-Editing-FT (Gangadhar and Stratos, 2024). These approaches facilitate the injection of multiple factual relations (up to the scale of around 10,000 factual relations) into the model and can be adapted for unlearning knowledge in LLMs by altering factual statements to end-of-sequence tokens – for example, changing "The mother tongue of David is French" to "The mother tongue of David is <|endoftext|>" (here “<|endoftext|>” is the end-of-sequence token in GPT-based models), effectively erasing specific information. While this strategy offers a potential pathway for knowledge unlearning, it may not surpass the effectiveness of our proposed method, which specifically focuses on the removal rather than the addition of factual relations. This distinction underscores the fundamental differences in approach between general model editing techniques and our targeted strategy for knowledge unlearning.
@@ 3.2 Previous Model Editing Strategy
As hypothesized and verified in Meng et al. (2022, 2023), the factual knowledge is mostly stored in the MLP layers, which leads to their strategy of updating the weight matrixes $W_{out}^{l}$ in Eq.(1). Meng et al. (2022) first identifies the layer in the model that contributes most to the related knowledge prediction, which we denote as $l_{0}$. Then the edit is performed on the parameter $W_{out}^{l_{0}}$. For simplicity, we denote $W_{0}$ as the specified parameter $W_{out}^{l_{0}}$ that needs to be updated. Inspired by Geva et al. (2020), the linear layer $W_{0}$ can act as key-value memories, associating input keys $K=\{k_{i}\}_{i=1}^{n}$ with corresponding values $V=\{v_{i}\}_{i=1}^{n}$. The following equation shows the relationship between $W_{0}$ and $K,V$:
@@ 4 Problem Setup
We define the problem Large Scale Knowledge Washing as: How to wash a certain large set of knowledge from the large language models without affecting the model’s reasoning ability? Here by washing the knowledge, we refer to the triplets that can be formed into single factual sentences. The knowledge set can be defined as follows:
@@ 5 Methodology
This formulation aims to disrupt the outputs significantly for inputs $K_{w}$, effectively "washing" the knowledge associated with $K_{w}$ from the model, thereby preventing accurate predictions based on $V_{w}$.
@@
Here $\beta$ is the hyperparameter used to control the tradeoff between the reasoning ability (related to $\parallel\hat{\Delta}K\parallel_{F}^{2}$) and the washing of previous knowledge (represented by $\parallel\hat{\Delta}K_{w}\parallel_{F}^{2}$). Then we simply set $\beta$ as 0.1 (an empirical value that should not affect the model’s ability on other tasks) and optimize the above objective to obtain the optimal $\Delta$. We visualize some details of the optimization in Figure 2.
@@ 5.1 Practical Consideration
5.2 Discussion of the Disentanglement of Knowledge and Reasoning
As demonstrated by Meng et al. (2022, 2023), the MLP (multi-layer perceptron) layers in transformers primarily store knowledge. However, our research also suggests that these layers significantly influence the model’s reasoning capabilities. This assertion is supported by experiments showing that modifications to the parameter $W_{0}$ can impact the model’s performance on reasoning tasks. Therefore, we propose that MLP layers are critical for both knowledge storage and reasoning processes. This paper explores strategies to disentangle these two functions by identifying alternative weight matrices that selectively diminish certain knowledge aspects while preserving, or minimally affecting, reasoning abilities. The possibilities of achieving this come from our hypothesis that knowledge storage and reasoning abilities can be separated within transformers. In this paper, we show the possibility of the disentanglement between knowledge and reasoning by washing a large amount of knowledge from the model while only minimally affecting the reasoning abilities.
@@ 6.1 Experimental Setup
To demonstrate the effectiveness of our method, we compare it with various knowledge editing and unlearning methods. The baselines for model-editing include: (1) FT: Simply finetune the model on the factual sentences formed from the triplets in $\mathcal{E}_{eos}$ in Eq.(6) (2) MEMIT (Meng et al., 2023): This state-of-the-art method can edit multiple layers of the model to perform thousands of edits simultaneously. (3) ME-FT (Model-Editing-FT) (Gangadhar and Stratos, 2024), which finetunes the model with only the loss on the span of $o_{i}$ in each sentence formed from $(s_{i},r_{i},o_{i})\in\mathcal{E}_{w}$. Irrelevant sentences are constructed as augmentations during training. For the knowledge unlearning category, the baselines are: (1) FT-UL (Finetune-Unlearning): Finetune the model on the sentences formed from the triplets $\mathcal{E}_{w}$ in Eq.(5) but with the reverse (i.e. negative) next-word-prediction loss function; (2) WOH (Who is Harry Potter) (Eldan and Russinovich, 2023): First train a reinforced model on the unlearning dataset, then update the target model to diverge from the reinforced model, based on the assumption that the reinforced model can better retain the unlearning dataset; (3) SeUL (Selective Forgetting) (Wang et al., 2024): Designates specific spans in the training data and uses a reversed next-word-prediction loss function on these spans for training.
@@
Consistent with previous studies (Meng et al., 2023; Gangadhar and Stratos, 2024), we employ GPT2-XL (1.5B parameters) and GPT-J-6B (6B parameters) as the backbone models for knowledge washing. The datasets used in our experiments are: (1) zsRE (Levy et al., 2017): A question-answering dataset with 19,086 facts. (2) CounterFactual (Meng et al., 2022): A dataset containing 21,929 counterfactual facts. After removing conflicting facts (Meng et al., 2023), 20,877 facts remain. (3) To facilitate large-scale knowledge washing, we utilize the latest Wikipedia dump, processing the relations following the guidelines provided in the repository [^2] https://github.com/neelguha/simple-wikidata-db. This results in 391,932,675 triplets. We then use gpt-3.5-turbo to rewrite each triplet into a sentence containing both the subject and the ground-truth answer. From 1,000,000 processed examples, we obtain 332,036 valid facts, creating the dataset referred to as Wiki-Latest.
@@
For evaluation, we employ two metrics to assess the extent of knowledge washing: (1) Accuracy: The model generates 10 tokens, and if the ground-truth answer is among the decoded output, it is considered a correct prediction. The accuracy is calculated as the percentage of correct predictions across the entire dataset. (2) QA-F1-Score: Using the metric from LongBench (Bai et al., 2023), we measure the F1-score between the generated output from the 10 tokens and the ground-truth answer. We measure the model’s reasoning ability with the library lm-evaluation-harness (Gao et al., 2023) on three tasks Lambda_openai (Radford et al., 2019; Paperno et al., 2016), HellaSwag (Zellers et al., 2019), and Arc_Easy (Clark et al., 2018). The descriptions of HellaSwag and Arc_Easy are shown in Appendix C.1. For most of the tables in the main paper, we only include the Accuracy and Perplexity from the evaluation results on Lambda_openai and leave the full table with the performances of the other two tasks in the appendix.
@@
GPT2-XL 1.0000 0.3704 0.5121 10.63 1.0000 0.2647 0.5121 10.63
FT 0.4208 0.2178 0.5275 9.72 0.1783 0.0930 0.5195 10.17
MEMIT 0.0462 0.0379 0.5156 10.52 0.1929 0.1439 0.4879 11.81
ME-FT 0.5091 0.2195 0.3881 21.95 0.1799 0.0878 0.3456 27.28
FT-UL 0.0000 0.0000 0.1126 $>10^{10}$ 0.0000 0.0000 0.0000 $>10^{10}$
WOH 0.5182 0.2017 0.5082 10.17 0.5978 0.1615 0.4619 13.15
SeUL 0.0957 0.0443 0.5108 10.66 0.0000 0.0000 0.2940 113.58
LaW 0.0050 0.0039 0.5108 10.86 0.1091 0.0905 0.4741 12.73
GPT-J-6B 1.0000 0.4043 0.6831 4.10 1.0000 0.4043 0.6831 4.10
FT 0.6181 0.2538 0.6887 4.02 0.3995 0.1646 0.6837 4.08
MEMIT 0.0553 0.0388 0.6815 4.14 0.2060 0.0759 0.6772 4.25
ME-FT 0.0751 0.0349 0.5178 8.53 0.2139 0.1183 0.4071 11.15
FT-UL 0.0000 0.0000 0.0000 $>10^{10}$ 0.0000 0.0000 0.0000 $>10^{10}$
WOH 0.6930 0.2829 0.6819 4.15 0.5396 0.1359 0.6833 4.19
SeUL 0.7422 0.3032 0.6815 4.15 0.5393 0.1395 0.6693 4.35
LaW 0.0454 0.0352 0.6701 4.35 0.0864 0.0334 0.6716 4.40

### v3

@@ Abstract
Large language models show impressive abilities in memorizing world knowledge, which leads to concerns regarding memorization of private information, toxic or sensitive knowledge, and copyrighted content. We introduce the problem of Large Scale Knowledge Washing, focusing on unlearning an extensive amount of factual knowledge. Previous unlearning methods usually define the reverse loss and update the model via backpropagation, which may affect the model’s fluency and reasoning ability or even destroy the model due to extensive training with the reverse loss. Existing works introduce additional data from downstream tasks to prevent the model from losing capabilities, which requires downstream task awareness. Controlling the tradeoff of unlearning existing knowledge while maintaining existing capabilities is also challenging. To this end, we propose LaW (Large Scale Washing), where we update the MLP layers in decoder-only large language models to perform knowledge washing, as inspired by model editing methods. We derive a new objective with the knowledge to be unlearned to update the weights of certain MLP layers. Experimental results demonstrate the effectiveness of LaW in forgetting target knowledge while maximally maintaining reasoning ability. The code is open-sourced at https://github.com/wangyu-ustc/LargeScaleWashing.
@@ 1 Introduction
Large Language Models (LLMs) are shown to memorize extensive knowledge or factual relations (Chen et al., 2022; Alivanistos et al., 2022; Youssef et al., 2023; Wang et al., 2024c). However, the memorization of knowledge in LLMs raises both moral and legal concerns. Factual knowledge may involve personal and sensitive information whose memorization can violate strict regulations (Legislature, 2018; Act, 1996; Parliament & of the European Union, 2016), and memorizing copyright content is also problematic – The New York Times [^1] https://nytco-assets.nytimes.com/2023/12/NYT_Complaint_Dec2023.pdf recently filed lawsuit against OpenAI to protect its copyright of articles. To prevent the undesired memorization of the above-mentioned knowledge, the simplest solution is perhaps to label data that has the potential to raise concerns in advance and exclude sensitive data from the pre-training stage. However, this solution needs exhaustive human effort and may not be feasible as the pretraining corpus for LLM is normally extremely large. This impossibility motivates the study of machine unlearning (Liu et al., 2024a; Yao et al., 2024; Si et al., 2023; Yao et al., 2023a; Zhang et al., 2023). When there are concerns about memorizing sensitive knowledge, these methods aim to update the LLM to forget that knowledge with a relatively small computational cost. Most of these methods are in the paradigm of defining an “unlearning” loss (essentially the reverse loss of Next-Word-Prediction on the unlearning dataset) and updating the full models by backpropagating from the loss. However, updating the model with backpropagation may hurt the model’s abilities in downstream tasks requiring reasoning. When the knowledge to be unlearned scales up, it may require extensive updates of the model parameters, which could even destroy the model (as shown in our experiments). Some efforts to overcome this limitation define a “utility” loss from specific downstream tasks and optimize both unlearning and utility losses (Liu et al., 2024a). However, the applications of these methods may be limited when we focus on the generalizability of LLMs where no downstream tasks are specified.
In this work, we focus on the Large Scale Knowledge Washing problem: How do we unlearn the knowledge at scale (termed knowledge washing) as cleanly as possible while minimizing the effects on the model’s reasoning ability? (as shown in Figure 1). We define reasoning ability as the model’s capacity to perform tasks that do not rely on prior domain knowledge, such as context-based question answering and mathematical reasoning. These tasks primarily require abstract pattern recognition and logical inference rather than memorized factual information. We hypothesize that the knowledge and reasoning abilities in LLMs are disentanglable, which gives rise to a feasible solution to the above problem. To address this, we design a novel method named LaW (Large Scale Washing), inspired by model editing techniques (Meng et al., 2022; 2023). Specifically, MEMIT (Meng et al., 2023) can perform extensive knowledge editing by identifying a subset of parameters in LLMs responsible for certain factual predictions and then modifying these parameters using a closed-form equation. Building on this concept, LaW first identifies the relevant subset of parameters and then formulates a new objective to update them in the context of knowledge washing. Unlike model-editing methods that aim to add factual relations to the model’s weights, LaW focuses on deleting factual relations. In the knowledge-washing scenario, while a closed-form solution similar to model editing is conceptually possible, practical constraints render some critical variables unavailable. Consequently, LaW introduces a novel objective that necessitates optimization rather than relying on a closed-form solution, incorporating several practical considerations to facilitate this process. Our primary contribution lies in demonstrating that LaW achieves superior and more thorough knowledge washing compared to existing methods. Meanwhile, our method is model-agnostic and applies to any transformer-based models with MLP layers. We evaluate LaW using two small-scale datasets and a newly created large-scale dataset derived from Wikipedia triplets, encompassing 332,036 facts. Experimental results reveal that LaW outperforms alternative approaches in effectively removing targeted knowledge, as evidenced by higher accuracy and QA-F1 scores on prompts derived from the triplets. Importantly, while LaW excels in unlearning, it maintains the model’s reasoning abilities to a reasonable extent, as validated through its performance on various reasoning tasks. This balance underscores LaW ’s effectiveness in achieving clean and comprehensive knowledge washing with minimal compromise on the model’s reasoning capabilities.
@@ 2 Related Work
Unlearning Knowledge in Large Language Model. Recent research has increasingly focused on the concept of machine unlearning in the context of large language models (LLMs), highlighting both its challenges and necessities (Liu et al., 2024a; Yao et al., 2024; Si et al., 2023; Yao et al., 2023a; Zhang et al., 2023; Wang et al., 2024b). Beyond addressing privacy concerns necessitating unlearning in LLMs, several studies have employed unlearning techniques to investigate the influence of specific subsets of training data on model performance (Isonuma & Titov, 2024; Zhao et al., 2024). To facilitate knowledge unlearning, various approaches have been proposed. One method involves retraining the LLM on the targeted dataset using a reverse loss function, coupled with training on an irrelevant dataset to preserve performance on unrelated tasks. This can be implemented through the addition of unlearning layers (Chen & Yang, 2023) or directly within the large language model itself (Eldan & Russinovich, 2023). Unlike these approaches, which apply to whole sequences in the unlearning subset, Wang et al. (2024) suggest focusing on specific spans within sequences to minimize disruption to unrelated tasks (Wang et al., 2024a). Furthermore, an alternative strategy known as in-context unlearning utilizes few-shot prompts to induce forgetting of specific datasets directly within the context of use, presenting a different approach from traditional training-based methods (Pawelczyk et al., 2023). In a distinct line of research, other methods target the mitigation of harmful outputs by collecting problematic prompts and applying techniques such as instruction tuning (Liu et al., 2024b) or reinforced learning (Lu et al., 2022) to prevent toxic responses.
@@
Model Editing of LLMs. Model editing in large language models pertains to the modification of factual relations within the models to integrate new world knowledge (Yao et al., 2023b). Initial approaches to model editing focused on single-fact adjustments, requiring the model to update one factual relation at a time. Prominent methods in this domain include ROME (Meng et al., 2022), MEND (Mitchell et al., 2022a), T-Patcher (Huang et al., 2023), and IKE (Zheng et al., 2023). These techniques, however, often face stability issues after multiple edits, complicating the process of batch editing, where multiple new factual relations are introduced simultaneously. In response to these challenges, advanced methods like GRACE (Hartvigsen et al., 2022) and SERAC (Mitchell et al., 2022b) have been developed for effective batch editing. Further advancements have tested these methodologies on larger models, such as GPT2-XL and GPT-J-6B, with techniques like MEMIT (Meng et al., 2023) and Model-Editing-FT (Gangadhar & Stratos, 2024). These approaches facilitate the injection of multiple factual relations (up to the scale of around 10,000 factual relations) into the model and can be adapted for unlearning knowledge in LLMs by altering factual statements to end-of-sequence tokens – for example, changing “The mother tongue of David is French” to “The mother tongue of David is <|endoftext|>” (here “<|endoftext|>” is the end-of-sequence token in GPT-based models), effectively erasing specific information. While this strategy offers a potential pathway for knowledge unlearning, it may not surpass the effectiveness of our proposed method, which specifically focuses on the removal rather than the addition of factual relations. This distinction underscores the fundamental differences in approach between general model editing techniques and our targeted strategy for knowledge unlearning.
@@ 3.2 Previous Model Editing Strategy
As hypothesized and verified in Meng et al. (2022; 2023), the factual knowledge is mostly stored in the MLP layers, which leads to their strategy of updating the weight matrixes $W_{out}^{l}$ in Eq.(1). Meng et al. (2022) first identifies the layer in the model that contributes most to the related knowledge prediction, which we denote as $l_{0}$. Then the edit is performed on the parameter $W_{out}^{l_{0}}$. For simplicity, we denote $W_{0}$ as the specified parameter $W_{out}^{l_{0}}$ that needs to be updated. Inspired by Geva et al. (2020), the linear layer $W_{0}$ can act as key-value memories, associating input keys $K=\{k_{i}\}_{i=1}^{n}$ with corresponding values $V=\{v_{i}\}_{i=1}^{n}$. The following equation shows the relationship between $W_{0}$ and $K,V$:
@@ 4 Problem Setup
We define the problem Large Scale Knowledge Washing as: How to wash a certain large set of knowledge from the large language models while minimizing the effects on the model’s reasoning ability? Here by washing the knowledge, we refer to the triplets that can be formed into single factual sentences. The knowledge set can be defined as follows:
@@ 5 Methodology
This formulation aims to disrupt the outputs significantly for inputs $K_{w}$, effectively ”washing” the knowledge associated with $K_{w}$ from the model, thereby preventing accurate predictions based on $V_{w}$.
@@
Here $\beta$ is the hyperparameter used to control the tradeoff between the reasoning ability (related to $\parallel\hat{\Delta}K\parallel_{F}^{2}$) and the washing of previous knowledge (represented by $\parallel\hat{\Delta}K_{w}\parallel_{F}^{2}$). Then we simply set $\beta$ as 0.1 (an empirical value that should not affect the model’s ability on other tasks) and optimize the above objective to obtain the optimal $\Delta$. We visualize some details of the optimization in Figure 2. We note that the regularization term $\frac{\|\hat{\Delta}K\|_{F}^{2}}{\|K\|_{F}^{2}}\leq\beta$ can be equivalently rewritten as $\frac{\hat{\Delta}(KK^{T})\hat{\Delta}}{KK^{T}}\leq\beta$, where $KK^{T}$ can be estimated using a WikiText dataset containing 100,000 text examples, as detailed in Appendix B.3 of Meng et al. (2023). Regarding $K_{w}$, it represents the input key to the target linear layer within the MLP that we aim to edit. Specifically, for the given knowledge statement "ChatGPT is developed by OpenAI", we input the prefix "ChatGPT is developed by" into the model and extract the hidden state corresponding to the last token "by" before it is fed into the target layer. This extracted hidden state serves as $K_{w}$ in practice.
@@ 5.1 Practical Consideration
5.2 Discussions
5.2.1 Disentanglement of Knowledge and Reasoning
As demonstrated by Meng et al. (2022; 2023), the MLP (multi-layer perceptron) layers in transformers primarily store knowledge. However, our research also suggests that these layers significantly influence the model’s reasoning capabilities. This assertion is supported by experiments showing that modifications to the parameter $W_{0}$ can impact the model’s performance on reasoning tasks. Therefore, we propose that MLP layers are critical for both knowledge storage and reasoning processes. This paper explores strategies to disentangle these two functions by identifying alternative weight matrices that selectively diminish certain knowledge aspects while preserving, or minimally affecting, reasoning abilities. The possibilities of achieving this come from our hypothesis that knowledge storage and reasoning abilities can be separated within transformers. In our experiments, we empirically demonstrate the feasibility of disentangling knowledge from reasoning by removing large amounts of knowledge without compromising reasoning abilities. Theoretically, we hypothesize that attention layers play a lesser role in encoding knowledge or reasoning, whereas MLP layers are more critical. Given that MLP layers decompose into keys and values, it is plausible that some are dedicated to reasoning, others to knowledge, and some to both. Investigating the extent to which specific keys and values contribute to reasoning could be worth future exploration.
5.2.2 Handling Output Behavior and Hallucination
Our algorithm does not aim to generate incorrect answers but rather disrupts the model’s confidence in the target knowledge, ensuring it “does not know” the correct answer. For example, given the input “ChatGPT is developed by,” the model would no longer confidently output “OpenAI.” If the model remains fluent, it may still generate a plausible continuation, such as another company name. However, since the erased knowledge is no longer accessible, the output becomes arbitrary or unrelated, reflecting the model’s updated state. While it may raise concerns such as hallucination, we argue this issue could potentially be mitigated through instruction tuning, where the model is explicitly trained to respond with “I don’t know” when lacking certain knowledge. In this paper, our focus is on removing specific knowledge learned during pre-training, whereas regulating the model’s response behavior falls under instruction tuning. This aligns with the observation in Zhou et al. (2023), where they claim that almost all knowledge is learned during pre-training, and instruction tuning could be used to adjust the models’ behaviors.
@@ 6.1 Experimental Setup
To demonstrate the effectiveness of our method, we compare it with various knowledge editing and unlearning methods. The baselines for model-editing include: (1) FT: Simply finetune the model on the factual sentences formed from the triplets in $\mathcal{E}_{eos}$ in Eq.(6) (2) MEMIT (Meng et al., 2023): This state-of-the-art method can edit multiple layers of the model to perform thousands of edits simultaneously. (3) ME-FT (Model-Editing-FT) (Gangadhar & Stratos, 2024), which finetunes the model with only the loss on the span of $o_{i}$ in each sentence formed from $(s_{i},r_{i},o_{i})\in\mathcal{E}_{w}$. Irrelevant sentences are constructed as augmentations during training. For the knowledge unlearning category, the baselines are: (1) FT-UL (Finetuning for Unlearning): Finetune the model on the sentences formed from the triplets $\mathcal{E}_{w}$ in Eq.(5) but with the reverse (i.e. negative) next-word-prediction loss function (i.e., the unlearning loss); (2) WOH (Who is Harry Potter) (Eldan & Russinovich, 2023): First train a reinforced model on the unlearning dataset, then update the target model to diverge from the reinforced model, based on the assumption that the reinforced model can better retain the unlearning dataset; (3) SeUL (Selective Forgetting) (Wang et al., 2024a): Designates specific spans in the training data and uses a reversed next-word-prediction loss function on these spans for training.
@@
Consistent with previous studies (Meng et al., 2023; Gangadhar & Stratos, 2024), we employ GPT2-XL (1.5B parameters) and GPT-J-6B (6B parameters) as the backbone models for knowledge washing. Please see the discussions of model choices in Appendix C.1. The datasets used in our experiments are: (1) zsRE (Levy et al., 2017): A question-answering dataset with 19,086 facts. (2) CounterFactual (Meng et al., 2022): A dataset containing 21,929 counterfactual facts. After removing conflicting facts (Meng et al., 2023), 20,877 facts remain. (3) To facilitate large-scale knowledge washing, we utilize the latest Wikipedia dump, processing the relations following the guidelines provided in the repository [^2] https://github.com/neelguha/simple-wikidata-db. This results in approximately 16,000,000 triplets. We then use gpt-3.5-turbo to rewrite each triplet into a sentence containing both the subject and the ground-truth answer. From 1,000,000 processed examples, we obtain 332,036 valid facts, creating the dataset referred to as Wiki-Latest.
@@
For evaluation, we employ two metrics to assess the extent of knowledge washing: (1) Accuracy: The model generates 10 tokens, and if the ground-truth answer is among the decoded output, it is considered a correct prediction. The accuracy is calculated as the percentage of correct predictions across the entire dataset. (2) QA-F1-Score: Using the metric from LongBench (Bai et al., 2023), we measure the F1-score between the generated output from the 10 tokens and the ground-truth answer. We measure the model’s reasoning ability with the library lm-evaluation-harness (Gao et al., 2023) on three tasks Lambda_openai (Radford et al., 2019; Paperno et al., 2016), HellaSwag (Zellers et al., 2019), and Arc_Easy (Clark et al., 2018). The descriptions of HellaSwag and Arc_Easy are shown in Appendix C.2. For the tables in the main paper, we report the average accuracy across three datasets Lambda_openai, arc_easy and hellaswag and leave the full table with all the other metrics in the appendix. Meanwhile, we add the results on GSM8k (Cobbe et al., 2021) in Appendix C.3.2. As the metrics of GSM8k are of a different scale, we do not include GSM8k when calculating the average accuracy in our main tables.
@@
GPT2-XL 1.0000 0.3704 0.5105 1.0000 0.2647 0.5105
FT 0.4208 0.2178 0.5049 0.1783 0.0930 0.5033
MEMIT 0.0462 0.0379 0.5130 0.1929 0.1439 0.4978
ME-FT 0.5091 0.2195 0.4801 0.1799 0.0878 0.3589
FT-UL 0.0000 0.0000 0.2398 0.0000 0.0000 0.1760
WOH 0.5182 0.2017 0.4993 0.5978 0.1615 0.4756
SeUL 0.0957 0.0443 0.4907 0.0000 0.0000 0.3558
LaW 0.0050 0.0039 0.5105 0.1091 0.0905 0.4890
GPT-J-6B 1.0000 0.4043 0.6560 1.0000 0.4043 0.6560
FT 0.6181 0.2538 0.6590 0.3995 0.1646 0.6544
MEMIT 0.0553 0.0388 0.6565 0.2060 0.0759 0.6502
ME-FT 0.0751 0.0349 0.5866 0.2139 0.1183 0.5112
FT-UL 0.0000 0.0000 0.1699 0.0000 0.0000 0.1707
WOH 0.6930 0.2829 0.6518 0.5396 0.1359 0.6535
SeUL 0.7422 0.3032 0.6514 0.5393 0.1395 0.6651
LaW 0.0000 0.0000 0.6468 0.0305 0.0125 0.6387